OpenStudy (anonymous):

find an equation of the mine containing (-3,2) and (3,1)

OpenStudy (anonymous):

y=(1/6)x+(1/2)

OpenStudy (anonymous):

line*

OpenStudy (anonymous):

above answer is not the right one or what?

OpenStudy (anonymous):

(1-2)/(3-(-3)) = -1/6 This is the gradient, therefore we can say y = -1/6x +c Choose a coordinate pair (3,1) and put into the above equation y =1 and x = 3 so 1= -1/6(3) + c therefore 1 = -1/2 +c so we can deduce from this that c = 3/2 by adding 1/2 to both sides. Now the equation says this y = -1/2x + 3/2 ..and that is the answer.